13. Artistic Adobes is considering growing its business by adding a paint machine that costs $100,000....

13.

Artistic Adobes is considering growing its business by adding a paint machine that costs

$100,000. The machine will generate an additional $33,800 in before-tax operating

income (excluding depreciation) for the next five years. At the end of five years, the

machine can be sold for $9,500. The machine falls into the MACRS 3-year class (pg. 184

or 205). Artistic’s marginal tax rate is 39%, and its required rate of return is 13%. Should

Artistic purchase the machine?
Answer #1

Artistic can purchase the machine as NPV is postive after taking in to account 13% cost of capital. Working is detailed below:

Calculation of Depreciation-Machinery
Year
1 2 3 4 5
Investment 100000 67000 22000 7000
Depreciation Rate 33% 45% 15% 7%
Depreciation 33000 45000 15000 7000
Value at Year End 67000 22000 7000 0
CASH FLOW CALCULATION OF THE PROJECT
Year
1 2 3 4 5
Operating Income before Tax (Excluding Depreciation)-A 33800 33800 33800 33800 33800
Depreciation-B 33000 45000 15000 7000 0
Sale of machinery-C 9500
Net Income Before Tax and After depreciation -D =(A-B+C) 800 -11200 18800 26800 43300
Tax @39%-E 312 -4368 7332 10452 16887
Post Tax Operaing Income-F= (D*E) 488 -6832 11468 16348 26413
Add: Depreciation 33000 45000 15000 7000 0
Cash Flow from Income Side -G=(F+B) 33488 38168 26468 23348 26413
Calculation of NPV Calculation of NPV Calculation of IRR
Cash Outflow Machiney 100000 Machinery         (100,000)
100000 0         (100,000)
Cash In Flow
Year Cash Inflow (A) Discounting Factor @13%-B Discounted Cash Flow-C=(A*B)
1 33488                0.8850                    29,635.40 1              33,488
2 38168                0.7831                    29,891.14 2              38,168
3 26468                0.6931                    18,343.65 3              26,468
4 23348                0.6133                    14,319.77 4              23,348
5 26413                0.5428                    14,335.92 5 26413
                 106,525.88
NPV of the Project is                       6,525.88 IRR 15.84%
As we can see there is positive NPV after taking in t acount the cost of 13%, Artistic Adobes can purchase the Machine